# Wireless Datasets for Aerial Networks

### v1

@@ XI UAV-Collected RSS Measurements for RF Source Localization
Figure 25: AFAR Challenge setup highlighting the UAV flight zone (blue) and the designated area where the UGV could be hidden (green). The UAV used during the experiment is shown in the upper-right corner [49].
@@ XI-A Description of Hardware and Software
The AFAR dataset includes data collected from both a DT simulation environment and a real-world wireless testbed, both deployed via the AERPAW platform. Each experimental run involved a UAV serving as the receiver and a UGV operating as the RF signal transmitter. The UAV and UGV were equipped with portable SDR nodes based on the USRP B205mini, capable of full-duplex operation over 70 MHz to 6 GHz.
Each SDR was connected to an Intel NUC 10 mini-PC equipped with an i7-10710U processor, 64 GB of RAM, and a 1 TB SSD, enabling real-time onboard signal processing. Transmissions employed a GNU Radio-based channel sounder that used a degree-12 Galois LFSR to generate a pseudo-random bit sequence (PRBS) of length 4095. The sequence was interpolated, pulse-shaped using a root-raised cosine filter, and transmitted at 2 MHz sampling rate over 3.0–4.2 GHz using a wideband antenna. On the receive side, frequency offset correction and correlation with the original PRBS enabled the extraction of CIR, from which RSS and received signal quality (RSQ) values were derived.

### v2

@@ XI UAV-collected RSS Measurements for RF Source Localization
Figure 24: AFAR Challenge setup highlighting the UAV flight zone (blue) and the designated area where the UGV could be hidden (green). The UAV used during the experiment is shown in Fig. 22.
@@ XI-A Description of Hardware and Software
The AFAR dataset includes data collected from both a DT simulation environment and a real-world wireless testbed, both deployed via the AERPAW platform. Each experimental run involved a UAV serving as the receiver and a UGV operating as the RF signal transmitter. The UAV and UGV were equipped with portable SDR nodes based on the USRP B205mini.
Each SDR was connected to an Intel NUC 10 mini-PC equipped with an i7-10710U processor, 64 GB of RAM, and a 1 TB SSD, enabling real-time onboard signal processing. Transmissions employed a GNU Radio-based channel sounder that generated a pseudo-random bit sequence (PRBS) of length 4095 using a degree-12 Galois LFSR. The sequence was pulse-shaped using a root-raised cosine filter and transmitted at a 2 MHz sampling rate over 3.0–4.2 GHz using a wideband antenna. At the receiver, frequency offset correction and correlation with the original PRBS enabled extraction of the CIR, from which RSS and received signal quality (RSQ) values were derived.